OpenStudy (anonymous):

Determine coefficients a and b such that p(x)=x^2+ax+b satisfies p(1)=2 and p′(1)=11.

OpenStudy (loser66):

where are you stuck?

OpenStudy (anonymous):

um ok so first you have to find the derivative right?

OpenStudy (loser66):

ok

OpenStudy (nincompoop):

maybe

OpenStudy (nincompoop):

you can solve for the p(x) without the derivative, you need it for the p'(x) however

OpenStudy (anonymous):

yeah i know so if p(1) = 2 then that means a+b is 1 right?

OpenStudy (nincompoop):

are you sure?

OpenStudy (anonymous):

i'm guessing that's wrong

OpenStudy (nincompoop):

\(p(1) = 2 = x^2+ax + b \rightarrow 2 = (1)^2+ a(1) + b \)

OpenStudy (anonymous):

yeah so if 1+a+b = 2 then a+b is 1

OpenStudy (nincompoop):

so a and b cannot be both 1

OpenStudy (nincompoop):

read the question "Determine coefficients a and b such..."

OpenStudy (anonymous):

oh no yeah i meant the sum of a+b is 1

OpenStudy (nincompoop):

now try to get the derivative of p(x)

OpenStudy (anonymous):

ok p'(x) is 2x + a

OpenStudy (nincompoop):

apply p'(1) = 11

OpenStudy (anonymous):

ok i got a = 9 and b= -8

OpenStudy (nincompoop):

aight

OpenStudy (anonymous):

thanks!

OpenStudy (nincompoop):

you seem to get this! laughing out loud

OpenStudy (nincompoop):

good job

OpenStudy (nincompoop):

your participation is greatly appreciated

OpenStudy (anonymous):

aww thank you for helping me :)

OpenStudy (nincompoop):

you helped yourself more than I helped you